A rectangle has a perimeter of 48, the ratio of the sides is 5:4. What is the length of the shortest side?
Simplest way to work this out, must be done only using scrap paper.

Found 2 solutions by Zucchini, rothauserc:
Answer by Zucchini(70)

To figure this out, you need to do 5x + 5x + 4x + 4x = 48. 18x = 48. X = 2 2/3. Now to find the shortest side multiply 4 by 2 2/3. 10 2/3 is the length of the shortest side.

Answer by rothauserc(4718)

the ratio of the sides is 5x:4x, then
2(5x) + 2(4x) = 48
10x +8x = 48
18x = 48
3x = 8
x = 8/3
then the shortest side is
4 * (8/3) = 32/3 = 10.666666667